# Frequency Decomposition-Driven Unsupervised Domain Adaptation for Remote Sensing Image Semantic Segmentation

### v1

@@ Abstract
[…]
1 Introduction
Driven by the rapid development of earth observation technology, semantic segmentation has become one of the most important tasks in practical remote sensing (RS) applications, such as land cover monitoring, land planning, and natural disaster assessment. In the era of deep learning, semantic segmentation models are normally trained by exploiting labeled datasets in a fully supervised manner (Wang et al., 2022b; Zhang et al., 2022b). However, their impressive performance is limited by the availability of high-quality annotated samples generated by the costly and laborious labeling process. Furthermore, they show poor generalization when the trained models are applied to data in other domains or unseen scenarios due to discrepancies between heterogeneous domains. To circumvent the domain shift, the Unsupervised Domain Adaption (UDA) approach has been investigated, which aims to transfer knowledge from the source and target domains and achieve the desired semantic segmentation performance in the target domain.
Figure 1: The main challenges of UDA semantic segmentation in remote sensing images. (a) Intra-domain: Ground objects exhibit varying scales and irregular boundaries. (b) Inter-domain: Larger-scale variations and more complicated boundaries are observed across domains.
Generally speaking, the core idea of UDA methods is to learn domain-invariant features across domains based on domain alignment, including discrepancy-based (Zhang et al., 2015), reconstruction-based (Ghifary et al., 2016) and adversarial-based (Ganin et al., 2016; Zhang et al., 2018; Chen et al., 2019) optimization principles. In particular, the adversarial-based methods were developed upon the generative adversarial network (GAN) that alternately trains a generator and a discriminator for domain alignment (Tsai et al., 2018; Vu et al., 2019). In semantic segmentation tasks, adversarial-based methods attempt to align two domains at image-level (Tasar et al., 2020; Li et al., 2022a), feature-level (Gao et al., 2023; Ma et al., 2023b) or output-level (Tsai et al., 2018; Vu et al., 2019; Ni et al., 2023). Since global context modeling is important for RS vision tasks, transformer (Vaswani et al., 2017), with its superiority by its ability to model long-range contextual information, has facilitated many GAN-based UDA methods by enhancing feature alignment (Li et al., 2022b; Zhang et al., 2022a; Ma et al., 2023b) or boosting the discriminative capability of the discriminator (He et al., 2022; Tu et al., 2022).
Notably, UDA on semantic segmentation of remote sensing images presents unique challenges. For example, the ground objects and their spatial relationships are complex in fine-resolution remote sensing images (Zheng et al., 2020). As shown in Fig. 1(a), ground objects such as buildings and trees vary widely in scale, even in the same location (intra-domain). Furthermore, the boundaries of ground objects are irregular, as illustrated in Fig. 1(b), making feature extraction very challenging. Finally, the spatial relationship between ground objects is more implicit. These challenges become more evident when images are acquired from different locations or domains (inter-domain), which incurs serious domain shift problems in UDA tasks. Therefore, we argue that both detailed local features and abstract global contexts are crucial for domain alignment in semantic segmentation. Although current methods perform well in aligning high-level features that mainly contain global contexts, the discrepancy in local details between the source and target domains is often neglected, leaving a gap in knowledge transfer across domains. Furthermore, it is still challenging to learn multiscale domain-invariant representations, which are beneficial for semantic segmentation tasks.
To address these problems, we propose a frequency decomposition-driven UDA method based on a global-local GAN model, namely FD-GLGAN, considering alignment in low-frequency global representations and high-frequency local information. Firstly, we propose a high/low-frequency decomposition (HLFD) technology that decomposes the feature map into high-frequency and low-frequency components through a CNN-based extractor and a window-based attention module, respectively. By aligning high- and low-frequency information simultaneously, the generator can learn more transferable features from the source domain. In particular, the plug-and-play HLFD can be easily embedded into other UDA backbones as a universal module to facilitate the learning of domain-invariant representations. Considering that the HLFD aligns the high- and low-frequency information at the encoding stage, to further exploit the decomposed information in adversarial learning strategy, a novel GAN model named GLGAN is proposed by exploiting an HLFD-based generator and a global-local discriminator (GLDis) both of which are derived from the global-local transformer block (GLTB) (Wang et al., 2022b). With the exploration of global contexts and local spatial information, the former facilitates the generation of accurate segmentation maps whereas the latter efficiently distinguishes the origins of segmentation maps. The contributions of this work are summarized in the following:
  • A novel plug-and-play HLFD module is proposed to learn high- and low-frequency domain-invariant features at multiple scales based on frequency-based feature decomposition. To the best of our knowledge, this is the first work that introduces the idea of frequency decomposition into UDA for remote sensing image semantic segmentation.
  • We develop a fully global-local adversarial learning-based UDA framework. It facilitates domain alignment by capturing cross-domain dependency relationships at different levels and exploiting global-local context modeling between two domains.
  • Extensive experiments on two benchmark datasets i.e., ISPRS Potsdam and Vaihingen, confirm that the proposed FD-GLGAN is more effective in overcoming the domain shift problem in cross-domain semantic segmentation than existing UDA methods.
The remaining parts of this paper are organized as follows. We first review the related works in Sec. 2 and the related theories in Sec. 3. Sec. 4 introduces the proposed HLFD and GLGAN in detail, whereas Sec. 5 elaborates on the experiment setup, experimental results and related analysis. Finally, the conclusion is given in Sec. 6.
2 Related Work
2.1 Cross-Domain Semantic Segmentation in Remote Sensing
Leveraging encoder-decoder-based networks (Ronneberger et al., 2015), numerous methods of semantic segmentation have emerged in the field of remote sensing (Kemker et al., 2018; Kotaridis and Lazaridou, 2021). Most are developed based on CNNs with residual learning and attention mechanisms to boost the representation learning (Diakogiannis et al., 2020; Kotaridis and Lazaridou, 2021; Li et al., 2021).To address the limitations of CNNs in modeling global representations, many transformer-based methods (Wang et al., 2022a; Ma et al., 2024) have been integrated under the encoder-decoder framework, leveraging the global context modeling capability of the self-attention mechanism (Vaswani et al., 2017; Liu et al., 2021).
In contrast to the straightforward semantic segmentation task, cross-domain semantic segmentation focuses on transferring knowledge from the source domain to the target domain. According to the transfer strategy, existing cross-domain semantic segmentation models can be divided into two main categories, namely the self-training approach (Yan et al., 2019; Liang et al., 2023) and the adversarial training approach (Tsai et al., 2018; Vu et al., 2019). The former generates pseudo-labels for the target domain using a model trained on the labeled source domain by exploiting pseudo-labels for the target domain. Alternatively, the latter approach focuses on producing semantic segmentation maps for the target domain that resemble those in the source domain. This is achieved by aligning features or outputs, which can be facilitated through the use of Generative Adversarial Networks (GANs).
Table 1: GAN-based Model classification according to the utilization strategies of global and local information.
Generator Discriminator Method
Local
Local
GANAI (Benjdira et al., 2019), TriADA (Yan et al., 2019), CAGAN (Liu et al., 2022), CCAGAN (Chen et al., 2022), MASNet (Zhu et al., 2023), MBATA-GAN (Ma et al., 2023b)
Global
Local
MuGCDA (Xi et al., 2023), ALE-UDA (Zhang et al., 2022a)
Global
Global
TransGAN (Jiang et al., 2021), ViTGAN (Lee et al., 2021), Swin-GAN (Wang et al., 2022d), HyperViTGAN (He et al., 2022), SWCGAN (Tu et al., 2022)
Global-Local
Global-Local
GLGAN (Ours)
2.2 GAN-based UDA in Remote Sensing
GAN-based UDA algorithms have been proven effective in cross-domain semantic segmentation for remote sensing imagery. Benjdira et al. (2019) proposed the first GAN-based model designed for cross-domain semantic segmentation from aerial imagery. On this basis, TriADA (Yan et al., 2019) introduced a triplet branch to improve the CNN-based feature extraction by simultaneously aligning features from the source and target domains. Recently, a self-training approach that employs high-confidence pseudo-labels was applied to incorporate category information (Liang et al., 2023; Ma et al., 2023a; Shen et al., 2024). Furthermore, SADA (Song et al., 2019) and CS-DDA (Zhang et al., 2020) developed subspace alignment methods to constitute a shared subspace with high-level features. To improve the high-level feature extraction, attention mechanisms were introduced into GAN-based UDA. More specifically, CAGAN (Liu et al., 2022) developed a covariance-based channel attention module to compute weighted feature maps, whereas CCAGAN (Chen et al., 2022) used a category-certainty attention module to align the category-level features. MASNet (Zhu et al., 2023) stored domain-invariant prototypical representations by employing a feature memory module while MBATA-GAN (Ma et al., 2023b) aligned high-level features by constructing a cross-attention-based transformer (Xu et al., 2021; Wang et al., 2022e). Compared with the methods above, our work attempts to fully explore global-local context modeling in the generator and discriminator, as shown in Table 1. Therefore, it can capture cross-domain global semantics and local detailed contexts while boosting the learning of domain-invariant features.
2.3 Decomposition-based UDA in Remote Sensing
The core idea of decomposition-based UDA is to split the features into different components before performing domain alignment in the corresponding subspaces. For instance, ToMF-B (Zhao et al., 2022) decomposed the cross-domain features into task-related features and task-irrelevant features by analyzing the gradients of the predicted score corresponding to the labels. Furthermore, ST-DASegNet (Zhao et al., 2024) introduced a domain disentangled module to extract cross-domain universal features and purify single-domain distinct features in a self-training framework. However, these techniques mainly concentrate on aligning high-level abstract features, overlooking the relationship between high/low-frequency details and the characteristics of ground objects. Therefore, it is crucial to address this shortfall by delving into frequency decomposition technology.
3 Preliminaries
Before introducing our proposed framework for UDA, we first revisit the basics of the multi-head self-attention (MHSA) design and the convolutional operation (Conv) from a frequency perspective. We employ the discrete Fourier transform (DFT) to analyze the feature maps regarding specific frequency components generated by MHSA and Conv. Let $X\in\mathbb{R}^{H\times W}$ and $Y\in\mathbb{R}^{C}$ represent an image in the spatial domain and its corresponding label, where $C$ denotes the number of classes. We transform $X$ into the frequency spectrum using the DFT $\mathcal{F}$: $\mathbb{R}^{H\times W}\rightarrow\mathbb{C}^{H\times W}$ and revert the signals of the image from frequency back to the spatial domain using the inverse DFT $\mathcal{F}^{-1}$: $\mathbb{C}^{H\times W}\rightarrow\mathbb{R}^{H\times W}$. For a mask $m\in\{0,1\}^{H\times W}$, the low-pass filtering $\mathcal{M}^{S}_{\mathrm{low}}$ and high-pass filtering $\mathcal{M}^{S}_{\mathrm{high}}$ with the filter size $S$ are formally defined as (Bai et al., 2022):
$\mathcal{M}^{S}_{\mathrm{low}}=\mathcal{F}^{-1}(m\odot\mathcal{F}(x)),\mathrm{where~{}}m_{i,j}=\begin{cases}1,~{}\mathrm{if~{}min}(|i-\frac{H}{2}|,|j-\frac{W}{2}|)\leq\frac{S}{2}\\ 0,~{}\mathrm{otherwise}\end{cases},$ (1)
and
$\mathcal{M}^{S}_{\mathrm{high}}=\mathcal{F}^{-1}(m\odot\mathcal{F}(x)),\mathrm{where~{}}m_{i,j}=\begin{cases}0,~{}\mathrm{if~{}min}(|i-\frac{H}{2}|,|j-\frac{W}{2}|)\leq\frac{\mathrm{min}(H,W)-S}{2}\\ 1,~{}\mathrm{otherwise}\end{cases},$ (2)
where $\odot$ stands for the Hadamard product, and $m_{i,j}$ denotes the value of $m$ at position $(i,j)$. According to the definitions in Eq. (1) and Eq. (2), the low-frequency components are shifted to the center of the frequency spectrum while the high-frequency components are shifted away from the center of the frequency spectrum.
Figure 2: The basic ideas of our frequency decomposition strategy. (a) Feature map extracted by the encoder, (b) feature map extracted further by MHSA and (c) its corresponding frequency spectrum; (d) feature map extracted further by Conv and (e) its corresponding frequency spectrum. The centroid represents low-frequency information, while the distance from it represents high-frequency information. More high-value points represent more information for the corresponding frequency.
Fig. 2 illustrates the contrasting behaviors that MHSA and Conv demonstrate. As shown in Fig. 2(b) and (d), MHSA aggregates feature maps, whereas Conv disperses them. Furthermore, as depicted in Fig. 2(c) and (e), Fourier analysis of feature maps reveals that MHSA focuses on low-frequency components, while Conv amplifies high-frequency components. In other words, MHSAs function as low-pass filters, while Conv serves as high-pass filters (Park and Kim, 2022). Consequently, MHSA and Conv complement each other and can be jointly employed for frequency decomposition. More details can be found in (Wang et al., 2020, 2022c).
Figure 3: The overview of the proposed FD-GLGAN that is comprised of a Generator, a Discriminator and the proposed HLFDs.
4 Proposed Method
In this section, we will first propose the overall structure of FD-GLGAN before elaborating on the proposed HLFD module as well as its generator and discriminator (GLDis). After that, the loss functions used in training will be presented.
Figure 4: The detailed structure of encoder and the detailed structure of the SwinTB.
4.1 Framework
Typical UDA tasks involve a labeled source dataset $D_{s}=\left\{(X_{s},Y_{s})\right\}^{n_{s}}$ and an unlabeled target dataset $D_{t}=\left\{X_{t}\right\}^{n_{t}}$ where $X_{s}$ represents an image in the source domain with its corresponding label $Y_{s}$ whereas $X_{t}$ denotes the unlabeled image in the target domain. Furthermore, $n_{s}$ and $n_{t}$ stand for the sample size of the source domain and the target domain, respectively. As the source domain and target domain possess different marginal and conditional distributions, i.e. $p_{s}(X_{s})\neq p_{t}(X_{t})$, the domain shift problem occurs. We assume that certain inherent similarities exist between different domains, which is valid as both domains share the same segmentation output space (Tsai et al., 2018; Wang et al., 2022e; Ma et al., 2023b).
As depicted in Fig. 3, the proposed FD-GLGAN is designed by capitalizing upon frequency decomposition and global-local context utilization. Specifically, a novel HLFD module is presented to align cross-domain representations by exploiting multiscale features generated by the encoder. Moreover, a novel global-local discriminator named GLDis further exploits the decomposed features to learn multiscale domain-invariant representations. In particular, the GLDis attempts to mimic the decoder to learn the domain-specific output space using the global-local representations.
4.2 Encoder
As illustrated in Fig. 4, the encoder comprises a Patch Partition layer and four successive stages. More specifically, the first stage contains one Embedding layer followed by the SwinTBs, whereas the following three stages are equipped with one Merging layer followed by SwinTBs each. We denote by $X_{s}\in\mathbb{R}^{{H\times W\times Z}}$ and $X_{t}\in\mathbb{R}^{{H\times W\times Z}}$ the training images from the source and target domains, respectively. Furthermore, $Z$ stands for the number of image channels, whereas $H$ and $W$ represent the height and width of the image, respectively. $X_{s}$ and $X_{t}$ are first split into non-overlapping patches of size $4\times 4$ by the Patch Partition layer. After that, the first-stage Embedding layer is applied to project these patches onto the encoding space before the SwinTBs gather image information on these patch tokens. As shown in Fig. 4, the SwinTB consists of four LayerNorm (LN) layers, two multilayer perceptrons (MLPs), one window-based multi-head self-attention (W-MHSA) and one shifted window-based multi-head self-attention (SW-MHSA) (Vaswani et al., 2017; Liu et al., 2021). After the first stage, three more successive stages are applied to generate the multiscale features denoted as $F_{s}^{\emph{1-4}}$ and $F_{t}^{\emph{1-4}}$ both of size ${\mathcal{S}}_{i}\times\frac{H}{2^{i+1}}\times\frac{W}{2^{i+1}}$ where ${\mathcal{S}}_{i}$ is the encoding dimension with $i\in\left\{1,2,3,4\right\}$ being the stage index. After that, these multiscale features are then fed into the corresponding HLFD module for frequency decomposition. Notably, the images from the source and the target domain are fed into the generator successively, that is, $F_{s}^{\emph{1-4}}$ and $F_{t}^{\emph{1-4}}$ are generated and collected, respectively.
Figure 5: The dataflow of the proposed multiscale HLFDs. The multiscale features from the source and target domains are decomposed by the corresponding scale HLFD into low-frequency and high-frequency decomposition features.
[…]
4.3 HLFD module
HLFD modules are employed to decompose the multiscale feature maps $F_{s}^{{1-4}}$ and $F_{t}^{{1-4}}$, as illustrated in Fig. 5. The feature map at each scale, including the features from the source domain and the target domain, will be decomposed into the high or low-frequency feature maps by the corresponding scale HLFD before these outputs are used to calculate the frequency alignment loss. Fig. 6 depicts the structure of the proposed HLFD module. Specifically, the HLFD module comprises two parallel branches, namely the low-frequency and high-frequency extractor. The former captures the low-frequency components by utilization of the W-MHSA (Liu et al., 2021). The following cross-shaped window context interaction (WCI) can maintain the global spatial consistency of ground objects. In contrast, the latter, designed for high-frequency components, has a structure of two parallel convolutional layers with kernel sizes of $3$ and $1$, respectively. Each convolutional layer is followed by a batch normalization operation. These branches facilitate the frequency decomposition of multiscale features based on the discussion in Sec. 3. The output low- and high-frequency components are then processed by an LN layer and an MLP, respectively. The final outputs of HLFD module denoted as $*LF_{s}^{\emph{*}}/LF_{t}^{\emph{*}}$ and $*HF_{s}^{\emph{*}}/HF_{t}^{\emph{*}}$ are used to computed the frequency alignment loss $L_{\rm{f}}$ that will be elaborated in Sec. 4.6.
Figure 7: (a) The detailed structure of GLTB. (b) The detailed structure of the decoder. (c) The detailed structure of the discriminator
4.4 Decoder
The decoder and GLDis that will be presented in Sec. 4.5 are both based on GLTB and designed to further exploit the decomposed features by gradually extracting global and local information. Their detailed structures and the GLTB are depicted in Fig. 7. Comparison of Fig. 6 and Fig. 7(a) suggests that the GLTB and HLFD modules share structural similarities. Specifically, the HLFD is replaced by global-local attention constituted of two parallel branches, namely the global and the local branches. The architecture of these branches closely mirrors that of the frequency extractor. However, in contrast to HLFD that is utilized for feature decomposition, the global-local attention leverages its two branches to investigate and incorporate global-local information. Hereby, the outputs from both the global and local branches are fused together before the resulting global–local context is further characterized by two convolution layers and a batch normalization operation.
@@ 4.4 Decoder
The high-level feature maps derived from the encoder are first fed into an individual GLTB whose output is processed by two stages of weighted sum operation and GLTB. In particular, GLTB is able to capture the global context and local details simultaneously. After that, the weighted sum operation is employed to extract domain-invariant features by adaptively fusing the residual features and those from the previous GLTB, as shown in Fig. 7(b). After the first three stages of the decoder, a weighted sum operation is used to fuse the residual connection from the first stage of the encoder and the deep global-local feature derived from previous GLTB before a classifier produces the final segmentation maps denoted by $P_{s}$ and $P_{t}$ of size $\mathbb{R}^{H\times W\times C}$ for the source and target domains, respectively.
@@
Finally, the supervised segmentation loss $L_{\rm{seg}}$ and the adversarial loss $L_{\rm{adv}}$ are also derived in the generator optimization stage. More specifically, $L_{\rm{seg}}$ is computed based on the source labels while $L_{\rm{adv}}$ is calculated using the adversarial strategy. More details about the loss functions will be elaborated in Sec. 4.6.
4.5 Discriminator
@@
Inspired by UNetformer (Wang et al., 2022b), we employ GLTB to exploit the semantic output space produced by the generator. As shown in Fig. 7(c), the proposed GLDis adopts a structure similar to the decoder but with inverted sampling operations. Specifically, it first down-samples the weighted self-information (WSI) distribution $I_{s}$ and $I_{t}$ derived from $P_{s}$ and $P_{t}$ before feeding the down-sampled data into a GLTB. More specifically, given the output segmentation maps of the source domain ${P}_{s}$ and the target domain ${P}_{t}$, their self-information are defined as $-\mathrm{log}{P}_{s}^{(H,W,c)}$ and $-\mathrm{log}{P}_{t}^{(H,W,c)}$, thereby the WSI maps ${I}_{s}$ and ${I}_{t}$ are computed by:
${I}_{s}^{(H,W,c)}=-{P}_{s}^{(H,W,c)}\cdot\mathrm{log}{P}_{s}^{(H,W,c)},$ (3)
${I}_{t}^{(H,W,c)}=-{P}_{t}^{(H,W,c)}\cdot\mathrm{log}{P}_{t}^{(H,W,c)}.$ (4)
@@
These information maps are regarded as the disentanglement of the Shannon Entropy. After the first GLTB, the resulting data is further processed by two successive stages of down-sampling followed by the GLTB, designed to gradually gather the global and local information in the segmentation map-generated features. Finally, the resulting data is down-sampled before being processed by a Classifier, which gives the predictions of the domain to which the input segmentation maps belong. Despite its appearance similar to the decoder structure, GLDis is tailored to learn domain-specific representations using an adversarial strategy. More specifically, the decoder upsamples the abstract semantic features, whereas GLDis down-samples the structured output space to identify the domain to which each input belongs. The final output of GLDis is a cross-entropy loss denoted by $L_{\rm{d}}$ in the discriminator training stage. More details about the loss function will be presented in Sec. 4.6.
4.6 Loss Functions

### v2

# Decomposition-based Unsupervised Domain Adaptation for Remote Sensing Image Semantic Segmentation

@@ III-C Encoder and Multiscale High/Low-Frequency Decomposition Modules
HLFD modules are employed to decompose the multiscale feature maps $\bm{F}_{s}^{\emph{1-4}}$ and $\bm{F}_{t}^{\emph{1-4}}$, as illustrated in Fig. 6. The feature map at each scale, including the features from the source domain and the target domain, will be decomposed into the high or low-frequency feature maps by the corresponding scale HLFD. After that, the final outputs of the HLFD modules denoted as $*\bm{LF}_{s}^{\emph{*}}/\bm{LF}_{t}^{\emph{*}}$ and $*\bm{HF}_{s}^{\emph{*}}/\bm{HF}_{t}^{\emph{*}}$ are used to compute the frequency alignment loss $L_{\rm{f}}$, which will be elaborated in Sec. III-F.
III-D Global-Local Decoder
@@
The decoder is developed based on GLTB [35] and designed to exploit the decomposed features by gradually extracting global and local information, as shown in Fig. 7(a). Specifically, the high-level feature maps derived from the encoder are first fed into an individual GLTB whose output is processed by two stages of weighted sum operation and GLTB. In particular, GLTB is able to capture the global context and local details simultaneously. After that, the weighted sum operation is employed to extract domain-invariant features by adaptively fusing the residual features and those from the previous GLTB. After the first three stages of the GLTB-based decoding, a weighted sum operation is used to fuse the residual connection from the first stage of the encoder and the deep global-local feature derived from previous GLTB before a classifier produces the final segmentation maps denoted by $\bm{P}_{s}$ and $\bm{P}_{t}$ of size $\mathbb{R}^{H\times W\times C}$ for the source and target domains, respectively, where $C$ stands for the number of ground object categories.
@@
Finally, the supervised segmentation loss $L_{\rm{seg}}$ and the adversarial loss $L_{\rm{adv}}$ are also derived in the generator optimization stage. More specifically, $L_{\rm{seg}}$ is computed based on the source labels while $L_{\rm{adv}}$ is calculated using the adversarial strategy. More details about the loss functions will be provided in Sec. III-F.
Figure 7: The detailed structures of (a) the decoder, and (b) the GLDis. They are similar in structure but different in function. The former is used for image decoding, while the latter is used for cross-domain information judgment.
III-E Global-Local Discriminator
@@
In the proposed discriminator, GLDis, we employ GLTB to exploit the semanftic output space produced by the generator. As shown in Fig. 7, the GLDis and the decoder are similar in structure but different in function. Specifically, it first down-samples the weighted self-information (WSI) distribution $\bm{I}_{s}$ and $\bm{I}_{t}$ derived from $\bm{P}_{s}$ and $\bm{P}_{t}$ before feeding the down-sampled data into a GLTB. Given the output segmentation maps of the source domain $\bm{P}_{s}$ and the target domain $\bm{P}_{t}$, the corresponding self-information is defined as $-\mathrm{log}\bm{P}_{s}$ and $-\mathrm{log}\bm{P}_{t}$, respectively. Consequently, the WSI maps $\bm{I}_{s}$ and $\bm{I}_{t}$ are calculated as follows:
$\bm{I}_{s}^{(h,w)}=-\bm{P}_{s}^{(h,w)}\cdot\mathrm{log}(\bm{P}_{s}^{(h,w)}),$ (5)
$\bm{I}_{t}^{(h,w)}=-\bm{P}_{t}^{(h,w)}\cdot\mathrm{log}(\bm{P}_{t}^{(h,w)}),$ (6)
@@
where the dimensions of $\bm{P}_{s}$, $\bm{P}_{t}$, $\bm{I}_{s}$ and $\bm{I}_{t}$ are $\mathbb{R}^{H\times W\times C}$, and $(h,w)$ represents the index position of one pixel. These information maps are regarded as the disentanglement of the Shannon Entropy. After the first GLTB, the resulting data is further processed by two successive stages of down-sampling, followed by the GLTB, which is designed to gradually gather the global and local information in the segmentation map-generated features. Finally, the resulting data is down-sampled and processed by a classifier, which predicts the domain to which the input segmentation maps belong. Despite its appearance similar to the decoder structure, GLDis is tailored to learn domain-specific representations using an adversarial strategy. More specifically, the decoder upsamples the abstract semantic features, whereas GLDis down-samples the structured output space to identify the domain to which each input belongs. The final output of GLDis is used to compute a cross-entropy loss denoted by $L_{\rm{d}}$ in the discriminator training stage. More details about the loss function will be presented in the next section.
III-F Loss Functions